# Stochastic Path Planning in Correlated Obstacle Fields

### v1

@@ 6.4 Offline Learning Framework Properties
    Monotonicity: if $X_{1}\subseteq X_{2}\subseteq X$, then $\mathbf{I}(\tilde{Y}_{X_{2}})\geq\mathbf{I}(\tilde{Y}_{X_{1}})$.
@@
    Submodularity: for all $X_{1}\subseteq X_{2}\subseteq X$ and any $x_{i}\notin X_{2}$,
    $\mathbf{I}(\tilde{Y}_{X_{1}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{1}})\;\geq\;\mathbf{I}(\tilde{Y}_{X_{2}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{2}}).$
@@
Normalization. $\mathbf{I}(\varnothing)=\tfrac{1}{2}\log\det(I)=0$.
Monotonicity. Let $X_{2}=X_{1}\cup\{x_{i}\}$. By the chain rule for mutual information,
$\mathbf{I}(\tilde{Y}_{X_{2}})-\mathbf{I}(\tilde{Y}_{X_{1}})=\text{MI}\big(Y_{x_{i}};\tilde{Y}_{x_{i}}\mid\tilde{Y}_{X_{1}}\big).$
For observations with Gaussian noise,
$\text{MI}\big(Y_{x_{i}};\tilde{Y}_{x_{i}}\mid\tilde{Y}_{X_{1}}\big)=\tfrac{1}{2}\log\Big(1+\tfrac{\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{1}})}{\sigma_{i}^{2}}\Big)\geq 0.$
So, $\mathbf{I}(\tilde{Y}_{X_{2}})\geq\mathbf{I}(\tilde{Y}_{X_{1}})$.
Submodularity. For $X_{1}\subseteq X_{2}$ and $x_{i}\notin X_{2}$,
$\mathbf{I}(\tilde{Y}_{X_{1}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{1}})=\tfrac{1}{2}\log\Big(1+\tfrac{\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{1}})}{\sigma_{i}^{2}}\Big).$
We can calculate $\mathbf{I}(\tilde{Y}_{X_{2}\cup\{x_{i}\}})-\mathbf{I}(\tilde{Y}_{X_{2}})$ in a similar way. In a Gaussian model, conditioning on more observations can only reduce conditional variance, $\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{2}})\leq\text{Var}(Y_{x_{i}}\mid\tilde{Y}_{X_{1}})$. Since $f(u)=\tfrac{1}{2}\log(1+u/\sigma_{i}^{2})$ is increasing, the marginal gain after observing $X_{1}$ is at least as large as after observing $X_{2}$, proving submodularity. ∎
Remark 6.3 (Justification for greedy sensing and disambiguation).
@@
Because $\textbf{I}(\cdot)$ is normalized, monotone, and submodular, the standard greedy selection achieves a (11/e)(1-1/e) approximation for maximizing I under a cardinality or budget constraint. This justifies our use of a simple greedy sensing and disambiguation strategy.
@@
Given state space and decision set are finite in our problem setting, assume every policy $\pi$ is proper, and the immediate cost is bounded $\sup_{s,d}|\mathcal{C}^{\scriptscriptstyle IG}(s,d)|<\infty$. Let $\bar{\bm{\rho}}$ be the mean transition probabilities derived from the posterior distribution over log-odds value $f(\mathbf{y}|\tilde{\rho}^{1},...,\tilde{\rho}^{t})$. The sequence $\{V_{i}\}$ generated by OPI framework with posterior sampling according to the update rule in Equation (2) converges almost surely to $V^{*}$ under transition information $\bar{\bm{\rho}}$.
@@
Define the Bellman operator under transition model $\bm{\rho}$ for a fixed policy $\pi$ as
$(\mathbf{T}_{\pi}^{\bm{\rho}}V)(s)=\mathcal{C}^{\scriptscriptstyle IG}(s,d^{\pi}(s))+\mathbf{E}_{\bm{\rho}}\left[V(s^{\prime})\right],$
and the optimal Bellman operator as
$(\mathbf{T}^{\bm{\rho}}V)(s)=\min_{d}\left\{\mathcal{C}^{\scriptscriptstyle IG}(s,d(s))+\mathbf{E}_{\bm{\rho}}\left[V(s^{\prime})\right]\right\}.$
Let the update rule be $V_{i}(s)=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$, where $\alpha_{i}=\frac{1}{N_{i}(s)}$, satisfying $\sum\alpha_{i}=\infty$ and $\sum\alpha_{i}^{2}<\infty$. $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ can be decomposed as $V^{\pi_{i},\bm{\rho}^{i}}+\omega_{i}$, where $V^{\pi_{i},\bm{\rho}^{i}}=\mathbf{T}_{\pi_{i}}^{\bm{\rho}^{i}}V_{i-1}$, and $\omega_{i}$ is the zero mean noise. Rewriting this update rule as
$V_{i}(s)=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}V^{\pi_{i},\bm{\rho}^{i}}+\alpha_{i}\omega_{i}.$
Define the posterior sampling noise as
$\eta_{i}=\mathbf{T}_{\pi_{i}}^{\bm{\rho}^{i}}V_{i-1}-\mathbf{T}_{\pi_{i}}^{\bar{\bm{\rho}}}V_{i-1},$
which satisfies $\mathbf{E}\left[\eta_{i}\right]=0$ due to the linearity of expectation. Following Lemma 2.5 in Chen, (2018), we have:
$\mathbf{T}_{\pi_{i}}^{\bar{\bm{\rho}}}V_{i-1}\leq\mathbf{T}^{\bar{\bm{\rho}}}V_{i-1}+\bm{\epsilon}.$
Since $\mathbf{E}\left[\omega_{i}\right]=0$, the inequality is established:
$V_{i}(s)$ $=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}V^{\pi_{i},\bm{\rho}^{i}}+\alpha_{i}\omega_{i}$
$\leq(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\bigl[\mathbf{T}^{\bar{\bm{\rho}}}V_{i-1}+\bm{\epsilon}\bigr]+\alpha_{i}(\eta_{i}+\omega_{i})$
$=(1-\alpha_{i})V_{i-1}(s)+\alpha_{i}\,\mathbf{T}^{\bar{\bm{\rho}}}V_{i-1}+\alpha_{i}\,\bm{\epsilon}+\alpha_{i}\,\tilde{\omega}_{i},$
where $\tilde{\omega}_{i}(s)$ is the combined zero-mean noise term. Then following the arguments in the proof of Theorem 2.1 in Chen, (2018) and Proposition 1 in Tsitsiklis, (2002), we have
$\lim_{i\rightarrow\infty}V_{i}=V^{*}\;\text{almost surely}.$
@@ 8.2 Competing Policies
These approaches use simulation to evaluate future trajectories, but differ in the base policy strategies and assumptions compared to our two-stage learning framework. While we learn high-quality base policy offline and continuously adapt them to evolving beliefs, these baselines rely on simple heuristic assumptions that remain static throughout traversal. We consider two approaches (Eyerich et al.,, 2010; Hou and Srinivasa,, 2022): (i) hindsight policy using rollout with a base policy that assumes perfect information about sampled environment during simulation, serving as a powerful benchmark for comparison in the literature, and (ii) optimistic rollout policy assuming all ambiguous obstacles in sampled environment are traverasable (i.e., optimistic assumption) during simulation, encouraging exploration in uncertain regions.
@@ 8.3 Illustrative Example
We use one example traversal region including 40 potentially blocked disks to illustrate how different policies and correlation modeling impact decisions and traversal cost.
Figure 5(a) presents the traversal from our proposed two-stage policy learning framework. The path takes a short, direct route after the information collection through sensing and disambiguation in the region near the starting vertex, reaching the goal with total cost 46.21. Figure 5(b)-5(c) include the traversal results using the baseline, DT policy, under two modeling assumptions, correlation-aware and independent. Compared with our two-stage strategy, the penalty policy is conservative and leads to much higher traversal cost. With correlation modeling, DT policy treats the central clustered region as high risk and tends to take detour around the region, resulting in a route with cost 90.36. The performance is worse when correlation is ignored. The policy cannot refine beliefs on statuses of nearby obstacles, repeatedly disambiguates locally with additional costs, increasing the cost to 120.43.
(a) Two-Stage Policy Learning
(b) Penalty Policy (correlation)
(c) Penalty Policy (independent)
Figure 5: Illustration using region with 40 potentially blocked disks under different policies and correlation assumptions.

### v2

@@ 6.4.2 Submodularity of Information Gain
    Monotonicity: If $A\subseteq B$, then $\mathbf{I}(A)\leq\mathbf{I}(B)$.
@@
    Submodularity (diminishing returns): For all $A\subseteq B\subseteq X$ and $i\in X\setminus B$,
    $\mathbf{I}(A\cup\{i\})-\mathbf{I}(A)\ \geq\ \mathbf{I}(B\cup\{i\})-\mathbf{I}(B).$
@@
(i) Normalization. $\mathbf{I}(\varnothing)=\tfrac{1}{2}\log\det(I)=0$.
(ii) Monotonicity. By the chain rule and independence of sensor noise across indices,
$\mathbf{I}(B)-\mathbf{I}(A)=\mathrm{MI}(Y_{B\setminus A};\,\widetilde{Y}_{B\setminus A}\mid\widetilde{Y}_{A})\ \geq 0.$
For a singleton addition $i\notin A$ this becomes
$\mathbf{I}(A\cup\{i\})-\mathbf{I}(A)=\tfrac{1}{2}\log\!\Big(1+\frac{s_{i\mid A}}{\sigma_{i}^{2}}\Big)\ \geq 0,$
where, by Gaussian conditioning (i.e., the Gaussian posterior variance of $Y_{i}$ given $\widetilde{Y}_{A}$ is),
$s_{i\mid A}:=\mathrm{Var}(Y_{i}\mid\widetilde{Y}_{A})=k_{ii}-k_{iA}(K_{AA}+\Sigma_{A})^{-1}k_{Ai}\ \in[0,k_{ii}].$
(iii) Submodularity. Let $A\subseteq B$ and $i\notin B$. Using the singleton marginal form above,
$\mathbf{I}(A\cup\{i\})-\mathbf{I}(A)=\tfrac{1}{2}\log\!\Big(1+\frac{s_{i\mid A}}{\sigma_{i}^{2}}\Big),\qquad\mathbf{I}(B\cup\{i\})-\mathbf{I}(B)=\tfrac{1}{2}\log\!\Big(1+\frac{s_{i\mid B}}{\sigma_{i}^{2}}\Big).$
It therefore suffices to show $s_{i\mid A}\geq s_{i\mid B}$. Write $C:=B\setminus A$ and apply block Gaussian conditioning (Schur complements):
$s_{i\mid A}-s_{i\mid B}=\ k_{iC\mid A}\,\bigl(K_{CC\mid A}+\Sigma_{C}\bigr)^{-1}\,k_{Ci\mid A}\ \ \geq\ 0,$
where $K_{CC\mid A}:=K_{CC}-K_{CA}(K_{AA}+\Sigma_{A})^{-1}K_{AC}$ and $k_{iC\mid A}:=k_{iC}-k_{iA}(K_{AA}+\Sigma_{A})^{-1}K_{AC}$. Hence $s_{i\mid B}\leq s_{i\mid A}$. Since $u\mapsto\tfrac{1}{2}\log(1+u/\sigma_{i}^{2})$ is increasing and concave on $0,\infty)$, the marginal gain decreases with the context, proving submodularity. ∎
Remark 6.4 (Noise models).
The proposition assumes independent per-index sensor noise (diagonal $\Sigma_{A}$), which is the sensing model used in the paper. If a fixed, subset-independent noise covariance $\Sigma_{0}\succ 0$ applies to all measurements simultaneously via a selection operator (observations $S_{A}Y+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\Sigma_{0})$), one can pre-whiten by $\Sigma_{0}^{-1/2}$ and derive an equivalent form with identity noise. In that case, a submodularity proof can be carried out analogously. For arbitrary subset-dependent correlated noise blocks $\Sigma_{A}$ that are not induced by restricting a single global $\Sigma_{0}$, submodularity of $\mathbf{I}$ need not hold in general, and extra care is needed.
@@
Because $\mathbf{I}$ is normalized, monotone, and submodular, the standard greedy selection achieves a (11/e)(1-1/e) approximation for maximizing $\mathbf{I}$ under a cardinality (or budget) constraint; see, e.g., (Nemhauser et al.,, 1978; Fisher et al.,, 1978) and subsequent sensor selection results.
6.4.3 Convergence under Posterior Sampling
We consider a finite MDP with state space $\mathcal{S}$, feasible decision set $\mathcal{D}(s)$, and bounded one–stage costs $\sup_{s,d}|\mathcal{C}(s,d)|<\infty$. Let $P_{\rho}(\cdot\mid s,d)$ denote the transition kernel parameterized by $\rho$. For a bounded $V:\mathcal{S}\to\mathbb{R}$, define the optimal Bellman operator under model $\rho$ by
$(\mathbf{T}^{\rho}V)(s)\;:=\;\min_{d\in\mathcal{D}(s)}\Big\{\mathcal{C}(s,d)+\sum_{s^{\prime}}P_{\rho}(s^{\prime}\!\mid s,d)\,V(s^{\prime})\Big\}.$
Assume one of the following settings holds:
  1. (S1)
    Discounted case: A discount $\gamma\in(0,1)$ is incorporated in $\mathcal{C}$ (or in the Bellman operator), so $\mathbf{T}^{\rho}$ is a $\gamma$–contraction in $\|\cdot\|_{\infty}$ with a unique fixed point $V^{\!*}(\rho)$.
  2. (S2)
    Stochastic shortest path (SSP): There exists an absorbing goal state $g$ with zero cost; all policies considered are proper (hit $g$ a.s. from any start), and the standard SSP conditions hold so that $\mathbf{T}^{\rho}$ has a minimal nonnegative fixed point $V^{\!*}(\rho)$ and value iteration converges to it.
Let $\{\mathcal{F}_{i}\}$ be the natural filtration generated by all randomness up to iteration $i$. At iteration $i$ we:
  • form the posterior and its mean $\bar{\rho}_{i}:=\mathbb{E}[\rho_{i}\mid\mathcal{F}_{i-1}]$;
  • compute a greedy policy $\pi_{i}(s)\in\arg\min_{d}\{\mathcal{C}(s,d)+\sum_{s^{\prime}}P_{\bar{\rho}_{i}}(s^{\prime}\!\mid s,d)\,V_{i-1}(s^{\prime})\}$;
  • sample a dynamics parameter $\rho_{i}$ from the current posterior (posterior sampling);
  • form a sampled policy-evaluation target $\widehat{\mathbf{T}}_{i}V_{i-1}(s)$, which is an unbiased estimator of $(\mathbf{T}^{\rho_{i}}_{\pi_{i}}V_{i-1})(s)$ (e.g., draw $S^{\prime}\!\sim P_{\rho_{i}}(\cdot\mid s,\pi_{i}(s))$ and use $\mathcal{C}(s,\pi_{i}(s))+V_{i-1}(S^{\prime})$).
We update per state with stepsizes $\alpha_{i}(s)\in(0,1]$:
$V_{i}(s)\ :=\ (1-\alpha_{i}(s))\,V_{i-1}(s)\ +\ \alpha_{i}(s)\,\widehat{\mathbf{T}}_{i}V_{i-1}(s).$
Assume (Robbins–Monro) for each $s$: $\sum_{i}\alpha_{i}(s)=\infty$, $\sum_{i}\alpha_{i}^{2}(s)<\infty$, and every state is updated infinitely often.
Let $\bar{\rho}_{i}:=\mathbb{E}[\rho_{i}\mid\mathcal{F}_{i-1}]$ denote the posterior mean at iteration $i$ and $\bar{\rho}$ a limit parameter (either the fixed mean if the dataset is frozen, or $\bar{\rho}_{i}\to\bar{\rho}$ a.s. if the posterior concentrates).
@@ 6.4.3 Convergence under Posterior Sampling
Under (S1) or (S2), and the conditions above, suppose in addition that:
  1. (A1)
    (Martingale-difference noise) The target noise
    $\xi_{i}(s)\ :=\ \widehat{\mathbf{T}}_{i}V_{i-1}(s)\ -\ (\mathbf{T}^{\rho_{i}}V_{i-1})(s)$
    satisfies $\mathbb{E}[\xi_{i}(s)\mid\mathcal{F}_{i-1}]=0$ and $\mathbb{E}[\xi_{i}(s)^{2}\mid\mathcal{F}_{i-1}]\leq C(1+\|V_{i-1}\|_{\infty}^{2})$ a.s. for some $C<\infty$.
  2. (A2)
    (Posterior sampling unbiasedness for fixed policy) The model mismatch term
    $\eta_{i}(s)\ :=\ (\mathbf{T}^{\rho_{i}}V_{i-1})(s)\ -\ (\mathbf{T}^{\bar{\rho}_{i}}V_{i-1})(s)$
    satisfies $\mathbb{E}[\eta_{i}(s)\mid\mathcal{F}_{i-1}]=0$ for all $s$.
  3. (A3)
    (Stable target) Either $\bar{\rho}_{i}\equiv\bar{\rho}$ for all $i$ (frozen posterior mean) or $\bar{\rho}_{i}\to\bar{\rho}$ a.s., and $\mathbf{T}^{\bar{\rho}_{i}}\to\mathbf{T}^{\bar{\rho}}$ uniformly on bounded sets.
Then $V_{i}\to V^{\!*}(\bar{\rho})$ almost surely as $i\to\infty$.
@@
Write the update as
$V_{i}(s)\ =\ (1-\alpha_{i}(s))\,V_{i-1}(s)\ +\ \alpha_{i}(s)\Big\{\ (\mathbf{T}^{\bar{\rho}}V_{i-1})(s)\ +\ \epsilon_{i}(s)\ +\ \zeta_{i}(s)\ \Big\},$
where we have decomposed the perturbation into
$\epsilon_{i}(s):=\underbrace{(\mathbf{T}^{\bar{\rho}_{i}}V_{i-1})(s)-(\mathbf{T}^{\bar{\rho}}V_{i-1})(s)}_{\to 0}\ +\ \underbrace{\big[(\mathbf{T}^{\bar{\rho}_{i}}_{\pi_{i}}V_{i-1})(s)-(\mathbf{T}^{\bar{\rho}_{i}}V_{i-1})(s)\big]}_{\leq 0},\qquad\zeta_{i}(s):=\eta_{i}(s)+\xi_{i}(s).$
By (A1)(A2), $\{\zeta_{i}(s),\mathcal{F}_{i}\}$ is a square-integrable martingale-difference sequence. By (A3), $\|\epsilon_{i}\|_{\infty}\to 0$ a.s. (the bracketed term is nonpositive and vanishes as $V_{i-1}\to V^{\!*}(\bar{\rho})$). Under (S1), $\mathbf{T}^{\bar{\rho}}$ is a contraction; under (S2), the standard monotone SSP convergence applies. Stochastic approximation for asynchronous value iteration with martingale noise then yields $V_{i}\to V^{\!*}(\bar{\rho})$ a.s. ∎
(A1) holds for unbiased Monte Carlo targets with bounded second moments. (A2) holds because, conditional on $\mathcal{F}_{i-1}$, $\rho_{i}$ is sampled from a posterior with mean $\bar{\rho}_{i}$ and $\mathbf{T}^{\rho}_{\pi}$ is linear in $P_{\rho}$ for fixed $\pi$. If the posterior is frozen, (A3) is trivial; if it concentrates, it holds when $P_{\bar{\rho}_{i}}\to P_{\bar{\rho}}$. If one updates only a subset of states per iteration, require that every state is visited infinitely often and apply the same argument componentwise (asynchronous SA). If an information–gain shaped cost $\mathcal{C}^{\scriptscriptstyle IG}$ is used, replace $\mathcal{C}$ by $\mathcal{C}^{\scriptscriptstyle IG}$ throughout.
6.4.4 Stability of Support Refinement
The distribution of value function at state $s$, $V(s)$, is categorical with support $\mathcal{X}(V(s))=\{x_{i}\}_{i=1}^{k}$, where the probabilities have a Dirichlet distribution with parameters $\{\alpha_{i}\}_{i=1}^{k}$. The mean is $\mu=\mathbf{E}\left[V(s)\right]=\frac{\sum_{i=1}^{K}\alpha_{i}x_{i}}{\sum_{i=1}^{k}\alpha_{i}}.$
Given a new observed cumulative cost $\mathcal{C}^{\scriptscriptstyle IG}(s)$ which is not in support and $\delta>0$ defining a neighborhood, we implement support refinement strategy through value contract. Specifically, we locate consecutive support points $x_{j},x_{j+1}$ within $\delta$ of $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$. If these points have not been observed in previous trajectories, we replace them with $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ and merge their Dirichlet parameters. If no neighbors exist within $\delta$, we insert $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ as a new support point with parameter $\alpha=1$.
Lemma 6.6.
Given a new observed cumulative cost $\mathcal{C}^{\scriptscriptstyle IG}(s)$ and $\delta>0$ defining a neighborhood, the support refinement through value contraction leads to a bounded change in the mean that shrinks as the total count grows, introducing no consistent upward or downward bias.
Proof.
We consider two cases based on the neighborhood structure.
(i) Two neighbors within $\delta$ distance. After replacing $x_{j},x_{j+1}$ with $\mathcal{C}^{\scriptscriptstyle IG}(s)$, the corresponding Dirichlet parameter is updated as $\alpha_{j}+\alpha_{j+1}+1$. Define $\alpha_{\mathrm{total}}=\sum_{i=1}^{k}\alpha_{i}$ and $\bar{x}_{j,j+1}=\frac{\alpha_{j}x_{j}+\alpha_{j+1}x_{j+1}}{\alpha_{j}+\alpha_{j+1}}$. The new mean becomes
$\mu^{\prime}=\frac{\alpha_{\mathrm{total}}\mu-(\alpha_{j}x_{j}+\alpha_{j+1}x_{j+1})+(\alpha_{j}+\alpha_{j+1}+1)\mathcal{C}^{\scriptscriptstyle IG}(s)}{\alpha_{\mathrm{total}}+1}.$
The difference between $\mu$ and $\mu^{\prime}$ is
$\mu^{\prime}-\mu$ $=\frac{\mathcal{C}^{\scriptscriptstyle IG}(s)-\mu}{\alpha_{\mathrm{total}}+1}+\frac{(\alpha_{j}+\alpha_{j+1})(\mathcal{C}^{\scriptscriptstyle IG}(s)-\bar{x}_{j,j+1})}{\alpha_{\mathrm{total}}+1}.$
The first term shrinks as the total count grows. For the second term, since $x_{j}$ and $x_{j+1}$ are within $\delta$ of $\mathcal{C}^{\scriptscriptstyle IG}(s)$, we have $|\mathcal{C}^{\scriptscriptstyle IG}(s)-\bar{x}_{j,j+1}|\leq\delta$. Therefore,
$|\mu^{\prime}-\mu|\leq\frac{|\mathcal{C}^{\scriptscriptstyle IG}(s)-\mu|+(\alpha_{j}+\alpha_{j+1})\delta}{\alpha_{\mathrm{total}}+1},$
leading to a bounded change that decreases with increasing $\alpha_{\mathrm{total}}$.
(ii) No neighbors within $\delta$ distance. A new support point is added with $\alpha=1$, giving $\mu^{\prime}=\frac{\alpha_{\mathrm{total}}\mu+\mathcal{C}^{\scriptscriptstyle IG}(s)}{\alpha_{\mathrm{total}}+1},$ and thus
$\mu^{\prime}-\mu=\frac{\mathcal{C}^{\scriptscriptstyle IG}(s)-\mu}{\alpha_{\mathrm{total}}+1},$
which also shrinks as $\alpha_{\mathrm{total}}$ increases, introducing no consistent upward or downward bias. ∎
@@ 8.2 Competing Policies
These approaches use simulation to evaluate future trajectories, but differ in the base policy strategies and assumptions compared to our two-stage learning framework. While we learn high-quality base policy offline and continuously adapt them to evolving beliefs, these baselines rely on simple heuristic assumptions that remain static throughout traversal. We consider two approaches (Eyerich et al.,, 2010; Hou and Srinivasa,, 2022): (i) hindsight policy using rollout with a base policy that assumes perfect information about sampled environment during simulation, serving as a powerful benchmark for comparison in the literature, and (ii) optimistic rollout policy assuming all ambiguous obstacles in sampled environment are traversable (i.e., optimistic assumption) during simulation, encouraging exploration in uncertain regions.
@@ 8.3 Illustrative Example
We use one example traversal region including 40 potentially blocked disks to illustrate how correlation modeling and the information gain bonus impact decisions and traversal costs.
Figure 5(a) presents the traversal from our proposed two-stage policy learning framework. The path takes a short, direct route after the information collection through sensing and disambiguation in the region near the starting vertex, reaching the goal with total cost 46.21. 5(b) shows the result when the information gain bonus is removed from our two-stage learning framework. This leads to a more conservative decision since the policy does not consider the potential value of new information that might benefit future decisions, instead directly choosing an obstacle-free path of cost 57.63. Figure 5(c)-(d) include the traversal results using the baseline policy, DT policy, under two modeling assumptions, correlation-aware and independent. Compared with our two-stage strategy, the penalty policy is myopic even with correlation. It enters the central clustered region without adequate strategic planning, and must detour around the obstacles when additional information reveals the risk, resulting in a route with cost 90.36. The performance is worse when correlation is ignored. The policy cannot refine beliefs on statuses of nearby obstacles, repeatedly disambiguates locally with additional costs, increasing the cost to 120.43.
Refer to caption Refer to caption
(a) Two-Stage Policy Learning (b) Two-Stage Policy (no information gain bonus)
Refer to caption Refer to caption
(c) Penalty Policy (correlation) (d) Penalty Policy (independent)
Figure 5: Illustration of different policy approaches for a region with 40 potentially blocked disks under various correlation assumptions (disk background color shows the ground truth, red and gray disks present the actual threats and false alarms, respectively).